# Photon Statistics for Fock and Coherent States Interfering in a Beamsplitter

### v1

@@ 1 Introduction
Over the past decades, numerous experiments have explored the behavior of single-photons, multi-photon Fock states, coherent states, and their superpositions passing through BSs [13, 14]. These studies have revealed rich photon statistics, nonclassical correlations, and phase space interference patterns. Notably, the observation of negative regions in the Wigner function serves as a hallmark of nonclassicality and has been confirmed in experiments involving Fock states and Schrödinger cat states [15, 16].
A key criterion to characterize the photon number statistics of a quantum state of light is Mandel’s $Q$ parameter [17], defined by the normalized variance of the photon number distribution relative to a Poisson distribution. This parameter quantifies deviations from Poissonian statistics: $Q=0$ corresponds to a Poissonian distribution (typical of coherent states), $Q>0$ indicates super-Poissonian (bunched) statistics typical of thermal light, and $Q<0$ reveals sub-Poissonian statistics, a signature of nonclassical light exhibiting photon antibunching [22]. Mandel’s $Q$ parameter thus provides a straightforward guideline to assess the quantum nature of light sources.
Closely related to $Q$ is the second-order coherence function $g^{(2)}(\tau)$, introduced by Glauber [1], which measures temporal correlations between photon detection events. At zero time delay, $g^{(2)}(0)$ can be expressed in terms of the photon number variance and mean, and thus relates directly to Mandel’s $Q$ parameter. This relation allows one to interpret the photon bunching or antibunching behavior of the field in terms of its number statistics, as we will see later.
@@ 1.1 Beamsplitter Theory
A bi-modal BS is a fundamental linear optical apparatus that mixes two input modes of the electromagnetic field into two output modes through a unitary transformation [23, 24]. In the quantum description, the BS acts on the bosonic creation operators $a_{a}^{\dagger}$ and $a_{b}^{\dagger}$ associated with the input modes $a$ and $b$, transforming them into output mode operators $a_{c}^{\dagger}$ and $a_{d}^{\dagger}$, as seen in Fig. (1).
Figure 1: Scheme of a lossless beamsplitter.
A common and widely used model for a lossless BS with complex amplitude reflection coefficient $r$ and transmission coefficient $t$, satisfying [23]
$|r|=|r^{\prime}|,\quad|t|=|t^{\prime}|,\quad|r|^{2}+|t|^{2}=1,$
$r^{*}t+r^{\prime}t^{*}=0,\quad r^{*}t+r^{\prime}t^{\prime*}=1,$
@@
where the factor $i$ in the reflection terms ensures the overall unitary operation and proper phase relations [25].
The inverse transformation reads:
@@
$a_{i},a_{j}^{\dagger}]=\delta_{ij},\quad i,j=a,b,c,d.$
@@
$U_{BS}(\theta,\phi)=\exp\left[\theta\left(e^{i\phi}a_{a}^{\dagger}a_{b}-e^{-i\phi}a_{a}a_{b}^{\dagger}\right)\right],$ (1.3)
with
  • $\theta\in[0,\pi/2]$ defines the mixing angle, with the amplitude transmission and reflection coefficients given by $|t|=\cos\theta$ and $|r|=\sin\theta$,
  • $\phi$ is a relative phase shift introduced between the reflected and transmitted fields.
@@ 1.2 Basic Photon Statistics
For a general state $\rho$, the mean photon number at each output is [26]:
@@
$g^{(2)}_{k}(0)==\frac{\langle\mathcal{N}_{k}^{2}\rangle-\langle\mathcal{N}_{k}\rangle}{\langle\mathcal{N}_{k}\rangle^{2}}.$ (1.8)
@@
$g^{(2)}_{k}(0)=1+\frac{Q_{k}}{\langle\mathcal{N}_{k}\rangle}.$ (1.9)
@@
The output state after passing through the BS is
@@
To obtain the explicit form of the output state in the photon-number basis, we expand each factor using the multi-binomial theorem:
@@ 2.1 Mean Photon Number
Note that the total photon number is conserved:
@@ 2.3 Mandel’s $Q$ Parameter
$Q_{c}=\frac{|t|^{2}|r|^{2}(n+m+2nm)}{n|t|^{2}+m|r|^{2}}-1,\ \text{for }n=m\neq 0,$ (2.17)
since vacuum is coherent, $Q_{c}=0$.
@@ 2.4 Second-order coherence function $g^{2}(0)$
Both parameters from equations (2.17) and (2.18) are plotted in figures (2) and (3), respectively.
Figure 2: Mandel $Q_{c}$ parameter for mode $c$ of a 50:50 BS with Fock-Fock state input of the type $|n\rangle_{a}|m\rangle_{b}$, where $|r|^{2}=|t|^{2}=1/2$. (a) Negative $Q_{c}$ values certify nonclassicality [22], demonstrating sub-Poissonian statistics from the input Fock states’ zero photon-number variance $(\Delta\mathcal{N})^{2}=0$. (b) At $m=n=1$, $Q_{c}=0$ despite the output being quantum (e.g., a HOM state), illustrating that $Q\geq 0$ does not guarantee classicality. (c) As $m\gg n$, $Q_{c}$ increases due to dominant super-Poissonian statistics. Numerical parameters $n\in\{0,1,2,5,10\}$ shows how input asymmetry affects quantum features.
Figure 3: $g^{(2)}(0)$ for mode $c$ of a 50:50 BS with Fock-Fock input of the type $|n\rangle_{a}|m\rangle_{b}$, where $|r|^{2}=|t|^{2}=1/2$. Notably, the Hong-Ou-Mandel (HOM) effect is observed for $n=m=1$, where perfect quantum interference between two indistinguishable photons leads to $g^{(2)}(0)=0$, indicating the complete suppression of simultaneous photon detections at the output.
@@ 2.5 Special Cases
When $n=1$ and $m=0$, equation (2.7) reduces to:
@@
which describes a single-photon in a superposition between the two output modes. Its statistics is:
$\langle\mathcal{N}_{c}\rangle=|t|^{2},\ (\Delta\mathcal{N}_{c})^{2}=|t|^{2}|r|^{2},\ Q_{c}=-|t|^{2},\ g^{2}_{c}(0)=0.$
2.5.2 One photon in each input $|1\rangle_{a}|1\rangle_{b}$
For $n=m=1$,
@@
In the special case of a 50:50 BS ($|t|=|r|=1/\sqrt{2}$), the term $|1\rangle_{c}|1\rangle_{d}$ vanishes due to destructive interference, leading to the well-known Hong-Ou-Mandel state:
@@
up to a phase. With photon statistics being:
$\langle\mathcal{N}_{c}\rangle=\langle\mathcal{N}_{d}\rangle$ $=1,\quad(\Delta\mathcal{N}_{c})^{2}=(\Delta\mathcal{N}_{d})^{2}=1,$
$\quad\ Q_{c}=Q_{d}$ $=0,\quad g^{2}_{c}(0)=g^{2}_{d}(0)=1.$
These results are well-known in quantum optics [23].
@@
We now consider a hybrid state where mode $a$ is a Fock state $|n\rangle$ and mode $b$ a coherent state $|\alpha\rangle$.
@@
where $\mathcal{D}_{b}(\alpha)=\exp(\alpha a_{b}^{\dagger}-\alpha^{*}a_{b})$ is the displacement operator for mode $b$.
The output state after the BS is:
$|out\rangle_{cd}$ $=U_{BS}\frac{(a_{a}^{\dagger})^{n}}{\sqrt{n!}}|0\rangle_{a}|\alpha\rangle_{b}$
$=\frac{(ta_{c}^{\dagger}+ira_{d}^{\dagger})^{n}}{\sqrt{n!}}\mathcal{D}_{c}(ir\alpha)\mathcal{D}_{d}(t\alpha)|0\rangle_{c}|0\rangle_{d}.$ (3.2)
Expanding the polynomial:
@@
The output state is:
@@ 3.1 Mean Photon Number
$\langle\mathcal{N}_{a}\rangle=\langle a_{a}^{\dagger}a_{a}\rangle=n,\quad\langle\mathcal{N}_{b}\rangle=\langle a_{b}^{\dagger}a_{b}\rangle=|\alpha|^{2}$ (3.5)
Thus:
@@ 3.3 Mandel $Q$ Parameter
Figure 4: Mandel $Q_{c}$ parameter for mode $c$ of a 50:50 BS with Fock-coherent state input of the type $|n\rangle_{a}|\alpha\rangle_{b}$, where $|r|^{2}=|t|^{2}=1/2$. (a) For $n=0$ (vacuum in mode $a$), $Q_{c}=0$ (Poissonian statistics) as the output remains coherent; (b) For $n\geq 1$, $Q_{c}$ exhibits sub-Poissonian statistics ($Q_{c}<0$) at small $\alpha$ due to Fock state antibunching, with $Q_{c}\rightarrow-1/2$ when $\alpha\rightarrow 0$, reflecting single-mode Fock state variance suppression; (c) As $\alpha$ increases, $Q_{c}>0$ (super-Poissonian) due to coherent state phase randomness. Numerical parameters: $\alpha\in[0,3]$, $n\in\{0,1,2,5,10\}$.
Figure 5: $g^{(2)}(0)$ for mode $c$ of a 50:50 BS with Fock-Fock input of the type $|n\rangle_{a}|m\rangle_{b}$, where $|r|^{2}=|t|^{2}=1/2$. As $|\alpha|\to 0$, the behavior approaches that of a single Fock state input, while for large $|\alpha|$, the statistics become increasingly classical.
@@ 3.4 Second-order coherence function $g^{2}(0)$
$Q$ and $g$ parameters are plotted in figures (4) and (5), respectively.

### v2

@@ 1 Introduction
Over the past decades, numerous experiments have explored the behavior of single-photons, multi-photon Fock states, coherent states, and their superpositions passing through BSs [13, 17, 14, 15, 16]. These studies have revealed rich photon statistics, nonclassical correlations, and phase space interference patterns. Notably, the observation of negative regions in the Wigner function [^1] The Wigner function provides a representation of a quantum quasiprobability distribution in phase space, a detailed analysis given in [20, 21]. serves as a hallmark of nonclassicality and has been confirmed in experiments involving Fock states and Schrödinger cat states [18, 19].
A key criterion to characterize the photon number statistics of a quantum state of light is Mandel’s $Q$ parameter [22], defined by the normalized variance of the photon number distribution relative to a Poisson distribution. Such distribution plays a central role in quantum optics as the reference case for photon counting statistics. It is a discrete probability distribution describing the likelihood of detecting $n$ photons when the mean photon number is $\langle\mathcal{N}\rangle$, and is written as
${\color[rgb]{0,0,0}P_{n}=\frac{\langle\mathcal{N}\rangle^{n}}{n!}e^{-\langle\mathcal{N}\rangle}.}$ (1.1)
Its defining feature is that the variance equals the mean, $\Delta\mathcal{N}^{2}=\langle\mathcal{N}\rangle$. Distributions broader than Poissonian ($\Delta\mathcal{N}^{2}>\langle\mathcal{N}\rangle$) are called super-Poissonian and correspond to photon bunching, where detection events tend to cluster together. In contrast, narrower distributions ($\Delta\mathcal{N}^{2}<\langle\mathcal{N}\rangle$) are called sub-Poissonian and correspond to photon antibunching, that is, the probability of detecting two photons in a very short time interval is smaller than for a Poissonian source, implying that photons tend to arrive one by one rather than in groups, a clear signature of nonclassical light [23].
Closely related to $Q$ is the second-order coherence function $g^{(2)}(\tau)$, introduced by Glauber [1], which measures temporal correlations between photon detection events. At zero time delay, $g^{(2)}(0)$ can be expressed in terms of the photon number variance and mean, and thus relates directly to Mandel’s $Q$ parameter. Experimentally, $g^{(2)}(\tau)$ is measured using setups such as the Hanbury Brown–Twiss (HBT) interferometer [24]. This relation allows one to interpret the photon bunching or antibunching behavior of the field in terms of its number statistics, as we will see later.
@@ 1.1 Beamsplitter Theory
A bi-modal BS is a fundamental linear optical apparatus that mixes input modes $a$ and $b$ of the electromagnetic field into output modes $c$ and $d$, through a unitary transformation [25, 26]. In the quantum description, the BS acts on the bosonic creation operators $a_{a}^{\dagger}$ and $a_{b}^{\dagger}$, transforming them into output mode operators $a_{c}^{\dagger}$ and $a_{d}^{\dagger}$, as seen in Fig. (1).
[…]
A common and widely used model for a lossless BS with complex amplitude reflection coefficient $r$ and transmission coefficient $t$, satisfying [25, 27]
$|r|=|r^{\prime}|,\quad|t|=|t^{\prime}|,\quad|r|^{2}+|t|^{2}=1,$ (1.2)
$^{*}t’+r’t^{*}=0,\quad r^{*}t+r’t’^{*}=0,$ (1.3)
@@
where the factor $i$ in the reflection terms ensures the overall unitary operation and proper phase relations [28]. The inverse transformation reads:
@@
$a_{i},a_{j}^{\dagger}]=\delta_{ij},\quad i,j=a,b,c,d.$ (1.6)
@@
$U_{BS}(\theta,\phi)=\exp\left[\theta\left(e^{i\phi}a_{a}^{\dagger}a_{b}-e^{-i\phi}a_{a}a_{b}^{\dagger}\right)\right].$ (1.7)
This operator is generated by bilinear combinations of bosonic operators, forming an $SU(2)$ algebra [29, 30, 31, 32]. The Baker–Campbell–Hausdorff (BCH) formula, together with equation (1.6), ensures that the linear transformation in equation (1.4) follows consistently from $U_{BS}$ [33].
The parameter $\theta\in[0,\pi/2]$ defines the mixing angle, which determines the amplitude transmission and reflection coefficients of the beamsplitter via
$|t|=\cos\theta,\quad|r|=\sin\theta.$
Physically, $\theta$ encodes energy conservation: the sum of reflected and transmitted intensities must equal the input. It also sets how strongly the two modes are mixed:
  • $\theta=0$ corresponds to perfect transmission ($t=1,r=0$), with no mixing of the input modes,
  • $\theta=\pi/4$ corresponds to a balanced 50:50 beamsplitter ($t=r=1/\sqrt{2}$), producing maximal superposition of the two input modes,
  • $\theta=\pi/2$ corresponds to perfect reflection ($t=0,r=1$), effectively swapping the modes.
The parameter $\phi$ is, mathematically, up to convention, however, it can also represent a physical phase shift, such as the one accumulated by one of the modes in an interferometric setup. For instance, in a Mach–Zehnder interferometer, a phase delay introduced in one arm appears as $\phi$, leading to observable interference fringes at the output [34, 35].
For realistic devices, however, losses render the BS transformation effectively non-unitary for the detected modes. These losses can be modeled by coupling each mode to vacuum ancillas, characterized by an overall efficiency $\eta$. The effect of such losses is to reduce the observed mean photon number $\langle\mathcal{N}\rangle$, drive the Mandel $Q$ parameter toward zero, and push the second-order coherence $g^{(2)}(0)$ toward unity. In Hong-Ou-Mandel (HOM) interference, symmetric losses preserve the dip depth but lower overall count rates, whereas asymmetric losses or background noise reduce the observed visibility [36, 37, 38].
@@ 1.2 Basic Photon Statistics
The mean photon number at each output is [39]:
@@
Here, $\rho=|out\rangle\langle out|$ is the density operator. The trace $\mathrm{Tr}(\rho\mathcal{O})$ denotes the sum of the diagonal elements of the operator $\mathcal{O}$ in the basis of $\rho$. In other words, it gives the quantum-mechanical average of the observable $\mathcal{O}$ for the system described by $\rho$ [23, 39].
@@
$g^{(2)}_{k}(0)=\frac{\langle\mathcal{N}_{k}^{2}\rangle-\langle\mathcal{N}_{k}\rangle}{\langle\mathcal{N}_{k}\rangle^{2}}.$ (1.12)
@@
$g^{(2)}_{k}(0)$ $=\frac{\langle\mathcal{N}_{k}^{2}\rangle-\langle\mathcal{N}_{k}\rangle}{\langle\mathcal{N}_{k}\rangle^{2}}$
$=1+\frac{\Delta\mathcal{N}_{k}^{2}-\langle\mathcal{N}_{k}\rangle}{\langle\mathcal{N}_{k}\rangle^{2}}$
$=1+\frac{Q_{k}}{\langle\mathcal{N}_{k}\rangle}.$ (1.13)
@@
Substituting input into output modes through equation (1.5), the state after passing through the BS becomes
@@
Equation (2.2) describes how a definite number of photons entering the two input ports are redistributed into the output modes $c$ and $d$. Each creation operator in the expansion represents the probability amplitude for a photon to be transmitted or reflected. To obtain the explicit form of the output state in the photon-number basis, we expand each factor using the multi-binomial theorem:
@@ 2.1 Mean Photon Number
For pure Fock state inputs, these reduce to $\langle\mathcal{N}_{c}\rangle=n$ and $\langle\mathcal{N}_{d}\rangle=m$ in the fully transmitting or reflecting limits. However, for mixed states, the mean photon numbers generally differ and must be calculated. Note that the total photon number is conserved:
@@ 2.3 Mandel’s $Q$ Parameter
$Q_{c}=\frac{|t|^{2}|r|^{2}(n+m+2nm)}{n|t|^{2}+m|r|^{2}}-1,\ \text{for }n=m\neq 0.$ (2.17)
For the vacuum state, $n=m=0$, the above equation is not directly applicable, as it would involve division by zero. Instead, $Q_{c}$ for the vacuum can be obtained by continuity, considering it as the $\alpha\to 0$ limit of a coherent state. In this limit, the vacuum is considered to have Poissonian statistics; This is a strictly quantum statement with no classical analogue: the vacuum possesses well-defined quantum fluctuations despite having zero photons on average, illustrating the inherently nonclassical nature of the vacuum [1].
@@ 2.4 Second-order coherence function $g^{2}(0)$
Both parameters from equations (2.17) and (2.18) are plotted in figures (2) and (3), respectively. As already mentioned, negative values of $Q_{c}$ indicate nonclassicality [23], reflecting the sub-Poissonian statistics of the input Fock states, for which the photon-number variance vanishes $(\Delta\mathcal{N})^{2}=0$. However, when $Q_{c}\geq 0$, no definite conclusion can be drawn. For instance, when $m=n=1$, the output is a Hong–Ou–Mandel state, yet $Q_{c}=0$. The Mandel parameter (and the single-mode $g$ function) captures only local photon-number fluctuations, but does not reveal correlations between modes (which, in the case of the HOM state, are highly quantum). In the limit $m\gg n$, $Q_{c}$ grows due to the dominance of super-Poissonian statistics.
This highlights an important distinction: while Mandel’s $Q$ and the single-mode $g^{(2)}(0)$ characterize local photon-number fluctuations, they are blind to inter-mode correlations. In practice, the hallmark of the HOM effect is not visible in $Q_{c}$ or $Q_{d}$, but rather in the cross-correlation $g^{(2)}_{cd}(0)$, which exhibits a pronounced dip due to the destructive two-photon interference. This shows that full characterization of quantum interference requires both local and nonlocal correlation functions [40].
Figure 2: Mandel $Q_{c}$ parameter for mode $c$ of a 50:50 BS with Fock-Fock state input of the type $|n\rangle_{a}|m\rangle_{b}$, where $|r|^{2}=|t|^{2}=1/2$.
Figure 3: $g^{(2)}(0)$ for mode $c$ of a 50:50 BS with Fock-Fock input of the type $|n\rangle_{a}|m\rangle_{b}$, where $|r|^{2}=|t|^{2}=1/2$.
@@ 2.5 Special Cases
For a single photon entering the beamsplitter ($n=1$, $m=0$), equation (2.7) reduces to
@@
which represents a photon in a coherent superposition of being in either output mode. It illustrates the fundamental principle of quantum interference at a beamsplitter: even a single photon exhibits wave-like behavior, with probability amplitudes split between the two outputs. Experimentally, this superposition can be observed in single-photon interference experiments using heralded photons from spontaneous parametric down-conversion [41].
The photon-number statistics in each output mode are:
$\langle\mathcal{N}_{c}\rangle=|t|^{2},\ (\Delta\mathcal{N}_{c})^{2}=|t|^{2}|r|^{2},\ Q_{c}=-|t|^{2},\ g^{(2)}_{c}(0)=0,$
indicating sub-Poissonian statistics and perfect antibunching, characteristic of single-photon states.
2.5.2 Two photons, one in each input $|1\rangle_{a}|1\rangle_{b}$
For the case $n=m=1$, the output state is
@@
For a symmetric 50:50 beamsplitter ($|t|=|r|=1/\sqrt{2}$), the $|1\rangle_{c}|1\rangle_{d}$ term vanishes due to quantum destructive interference between indistinguishable photons, giving rise to the HOM effect [11], up to a global phase. Such effect is routinely used to test photon indistinguishability in experiments with single-photon sources, e.g., from quantum dots [42]:
@@
The photon-number statistics for each output mode are
$\langle\mathcal{N}_{c}\rangle=\langle\mathcal{N}_{d}\rangle=1,\ (\Delta\mathcal{N}_{c})^{2}=(\Delta\mathcal{N}_{d})^{2}$ $=1,$
$Q_{c}=Q_{d}=0,\ g^{(2)}_{c}(0)=g^{(2)}_{d}(0)=1,$
Indicating Poissonian statistics for each mode individually, despite the strong quantum correlations between them.
@@
We now consider a hybrid state where mode $a$ is a Fock state $|n\rangle$ and mode $b$ a coherent state $|\alpha\rangle$:
@@
where $\mathcal{D}_{b}(\alpha)=\exp(\alpha a_{b}^{\dagger}-\alpha^{*}a_{b})$ is the displacement operator for mode $b$. Particularly, the displacement operator shifts the quantum state in phase space: it translates both the quadratures of the field without altering their quantum uncertainties [25]. In this context, a coherent state $|\alpha\rangle$ can be viewed as the vacuum state displaced by $\alpha$ in phase space, i.e., $\mathcal{D}(\alpha)|0\rangle=|\alpha\rangle.$
After the beamsplitter, the output state is
$|out\rangle_{cd}$ $=U_{BS}|in\rangle_{ab}$
$=U_{BS}\frac{(a_{a}^{\dagger})^{n}}{\sqrt{n!}}|0\rangle_{a}|\alpha\rangle_{b}$
$=\frac{(ta_{c}^{\dagger}+ira_{d}^{\dagger})^{n}}{\sqrt{n!}}\mathcal{D}_{c}(ir\alpha)\,\mathcal{D}_{d}(t\alpha)\,|0\rangle_{c}|0\rangle_{d},$ (3.2)
where we have used the beamsplitter transformations of equation (1.5) and the displacement property.
Now, let us expand the operator $(ta_{c}^{\dagger}+ira_{d}^{\dagger})^{n}$ using the binomial theorem:
@@
Substituting this expansion into the output state, we obtain:
@@ 3.1 Mean Photon Number
$\langle\mathcal{N}_{a}\rangle=\langle a_{a}^{\dagger}a_{a}\rangle=n,\quad\langle\mathcal{N}_{b}\rangle=\langle a_{b}^{\dagger}a_{b}\rangle=|\alpha|^{2}.$ (3.5)
The mean for the coherent state comes precisely from (one of) its definitions: the coherent state is an eigenstate of the annihilation operator:
$a|\alpha\rangle=\alpha|\alpha\rangle.$ (3.6)
This property implies shot-noise level: relative fluctuations scale as $1/\sqrt{\langle\mathcal{N}\rangle}$, and thus coherent states define the classical benchmark with $g^{(2)}(0)=1$, separating sub-Poissonian (antibunched) from super-Poissonian (bunched) light. Hence:
@@ 3.3 Mandel $Q$ Parameter
Figure 4: Mandel $Q_{c}$ parameter for mode $c$ of a 50:50 BS with Fock-coherent state input of the type $|n\rangle_{a}|\alpha\rangle_{b}$, where $|r|^{2}=|t|^{2}=1/2$. Numerical parameters: $\alpha\in[0,3]$, $n\in\{0,1,2,5,10\}$.
Figure 5: $g^{(2)}(0)$ for mode $c$ of a 50:50 BS with Fock-coherent input of the type $|n\rangle_{a}|\alpha\rangle_{b}$, where $|r|^{2}=|t|^{2}=1/2$.
@@ 3.4 Second-order coherence function $g^{2}(0)$
$Q$ and $g$ parameters are plotted in figures (4) and (5), respectively. For $n=0$ (vacuum in mode $a$), $Q_{c}=0$ (and $g^{2}_{c}(0)=1$), indicating Poissonian statistics. For $n\geq 1$, $Q_{c}$ becomes negative at small $\alpha$ due to Fock-state antibunching, approaching $Q_{c}\rightarrow-1/2$ as $\alpha\rightarrow 0$, which reflects the suppression of single-mode photon-number variance. In this regime, particularly for $n=1$ and $\alpha\to 0$, a single photon cannot split into two counts, leading to fluctuations below the Poisson level. As $\alpha$ increases, however, the coherent-state component dominates, introducing phase randomness and driving $Q_{c}$ toward positive values (super-Poissonian statistics). Equation (3.1) thus illustrates the continuous transition between quantum and classical regimes: for small $|\alpha|$, the Fock contribution governs the behavior, yielding antibunching and negative $Q$, whereas for large $|\alpha|$, the coherent field washes out the quantum correlations. This behavior highlights how realistic light sources naturally interpolate between ideal quantum and classical limits.